OpenStudy (anonymous):

Is x^(2)y' + e^(x)*y=4 a differential equation that is linear?

OpenStudy (across):

Yes, it is linear.

OpenStudy (anonymous):

it goes to: (dy/dx)x^2 + ye^(x) = 4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

and then you divide both sides by x^2?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

ahh i see. do i divide ye^(x) by x^2 as well?

OpenStudy (dumbcow):

yes then you have to find the integrating factor

OpenStudy (anonymous):

oh okay, i got it, thanks guys!

OpenStudy (unklerhaukus):

\[x^2y' + e^x y=4\] \[y' + \frac{e^x}{x^2} y=\frac{4}{x^2}\] This is linear because it is of the form; \[y'+p(x)y=q(x)\] Where \( p(x)= \frac{e^x}{x^2}\) \(q(x)=\frac{4}{x^2}\)